The Blair Company’s three assembly plants are located in California, Georgia, and New Jersey.  Previously, the company purchased a major subassembly, which becomes part of the final product, from an outside firm.  Blair has decided to manufacture the subassemblies within the company and must now consider whether to rent one centrally located facility (e.g. in Missouri, where all subassemblies would be manufactured) or to rent three separate facilities, each located near one of the assembly plants, where each facility would manufacture only the subassemblies needed for the nearby assembly plant.  A single, centrally located facility, with a production capacity of 18,000 units per year, would have fixed costs of $900,000 per year and variable cost of $250 per unit.  Three separate decentralized facilities, with production capacities of 8,000, 6,000 and 4,000 units per year, what have fixed costs of $475,000, $425,000 and $400,000, respectively, and variable costs of $225 per unit, owing primarily to the reduction in shipping costs.  The current production rates at the three assembly plants are 6,000, 4,500, and 3,000 units, respectively, to satisfy 13,500 units demanded.

  1. Assuming that the current production rates and demand are maintained at the three assembly plants, which alternative should management select?
  2. If demand for the final product were to increase to production capacity, which alternative would be more attractive?
  3. What additional information would be useful before making a decision?
Answer #1

The Blair Company has multiple plant locations.

a.

One centralized plant in Missouri:

TC = $900,000 + (6,000 + 4,500 + 3,000) $250 = $4,275,000

Three subassembly plants:

TC = ($475,000 + $425,000 + $400,000) + (6000 + 4500 + 3000) $225 = $4,337,500

One centralized plant would be cheaper.

b.

One centralized plant at full capacity:

TC = $900,000 + (18,000) $250 = $5,400,000

Three plants:   TC = ($475,000 + $425,000 + $400,000) + (8,000 + 6,000 + 4,000) $225

= $5,350,000.

Three plants should be built.

c.     

It would be helpful to have a credible forecast of the future demand for subassemblies.